OpenStudy (anonymous):

How do you convert the ellipse 9x^2 + 16y^2 = 144 into polar coordinates?

OpenStudy (anonymous):

x=r cos theta and y=r sin theta You can plug those in and mess with it to see if you can get something nice. There are also some versions used in astronomy. The include the definition of eccentricity... you would have to look those up to convert to those.

OpenStudy (anonymous):

Yeah I have plugged those in for the equation. What do I do with the 144 though? Is that used in solving for r? 144r^2 = 9x^2 + 16y^2, which results in r = x/4 + y/3 ?

OpenStudy (anonymous):

Once you do the substitution, the x's and y's should be gone. I don't think that you will find any solution that is elegant without going all the way to those astronomy versions.

OpenStudy (anonymous):

I am doing a double integral with C being that ellipse, then using Green's Theorem to convert into a double integral. I'm just having trouble translating the given ellipse into proper bounds of integration

OpenStudy (anonymous):

The standard form for the ellipse would actually be\[\frac{x ^{2}}{16}+\frac{y ^{2}}{9}=1\]

OpenStudy (anonymous):

Oh, wow... would parametric be better? That's easier.

OpenStudy (anonymous):

x=4cos(t) and y=3sin(t) over [0,2pi]

OpenStudy (anonymous):

Ok I will give that try, just need to refresh myself on parametric stuff to find the bounds of integration using parametric equations

OpenStudy (anonymous):

And I need to refresh on Green's to see what that takes... good luck.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (anonymous):

From your perspective, they can be seen as x and y components of vector valued functions. Substitute x values with 4 cos t, y values with 3 sin t, dx with -4 sin t, and dy with 3 cos t. Then integrate from 0 to 2pi... I think. :)

OpenStudy (anonymous):

Maybe I should just be quiet... I'm really rusty. :)

OpenStudy (anonymous):

What I ended up doing was using Green;s theorem to convert the equation in the original integral in M(x,y) and N(x,y). Then I did the partial derivative of M with respect to y and the partial derivative of N with respect to x

OpenStudy (anonymous):

Both of those are 4x, so given Green;s theorem, that would be 4x - 4x

OpenStudy (anonymous):

Resulting in an area of 0, because the integral of 4x-4x would be 0

OpenStudy (anonymous):

Closed path and conservative vector field should result in zero. Did you test if it was conservative to start with?

OpenStudy (anonymous):

Wait... yeah... that is the test... It's conservative!

OpenStudy (anonymous):

Yeah, so when using Green's theorem I guess if it is conservative, then the area will be 0

OpenStudy (anonymous):

Yes, because to use Green's requires a closed path.

OpenStudy (anonymous):

Exactly. But for future reference, it's best to use parametric equations for ellipses, circles or lines and polar coordinates for other curves?

OpenStudy (anonymous):

I would say so... parametric is almost the same as vector valued functions, so they relate well.

OpenStudy (anonymous):

Alright awesome. Calc 3 is almost done, hallelujah!

OpenStudy (anonymous):

Thanks again for all your help

OpenStudy (anonymous):

Right on! You're welcome.